If log_2x=7, evaluate log_2 16x^2
Please any help is appreciated thank you

Found 2 solutions by solver91311, london maths tutor:
Answer by solver91311(24713)





Since we know



means that



We can say



But we also know that



So



Which is to say:



Use the Power Rule

because



John


Answer by london maths tutor(243)

If log_2x=7, evaluate log_2 16x^2
Rewrite log2 (16x^2) as:
2[log2 (16x)]
=2(log2 16) + 2(log2 x) , log2 x = 7
=2(log2 2^4) + (2*7) , log2 2 = 1
=(2*4) + 14
= 8 + 14
= 22